# Periodic Korteweg-de Vries soliton potentials generate magnetic field strength with excellent quasisymmetry

### v3

@@ Appendix E Additional configurations that satisfy the integrated KdV equation (12)
The precise QA configuration with larger negative shear in Fig. 21 was obtained by a series of optimizations starting from the precise QA.Landreman and Paul (2022) Specifically, we start from the same objective function used to optimize the precise QA, and the precise QA configuration itself. The mean shear $\bar{\hat{s}}$ was extracted from the precise QA configuration, and a new least-squares term $(\bar{\hat{s}}-\bar{\hat{s}}_{*})^{2}$ was added to the objective function, where the target value for mean shear $\bar{\hat{s}}_{*}$ is taken to be slightly different from the extracted value. This new objective is then optimized locally, using the precise QA as the initial state. This ensures that the additional least-squares term is almost zero, so the precise QA is approximately an optimum to the modified objective function. Hence, the new local optimum can be expected to be close to the initial configuration. This process is then repeated starting from the slightly modified optimum, each step generating a similar equilibrium with a slightly different mean shear. Mean shear is here defined as $b/\bar{\iota}$, where $b$ is the result of fitting a first-order polynomial $a+bs$ to the iota profile, where $s$ is the normalized toroidal flux; $\bar{\iota}$ is the rotational transform averaged over all radii. The mean shear in the original precise QA is about $-0.018$, and we incremented the target negative shear in steps of $0.005$.
@@
We show equilibria that were not derived from the precise QA/QH in Figures 23 and 24. Instead, they were only optimized for quasisymmetry on the outermost flux surface, with the quasisymmetry degrading as the axis is approached. For both cases, the ratio of symmetry-breaking modes to quasisymmetric modes near the axis is $\sim 10^{-3}$. While this may seem a small value, the resulting ripple creates enough noise in the $(dB/d\ell)^{2}$ vs. $B$ plot to render it meaningless, and the error for all fits is $\sim 1$ in that region. The symmetry-breaking modes were filtered out when calculating the roots shown in panels (c). Note that the equilibrium shown in Fig. 23 is the same one used in Fig. 2 in the main text.

### v4

@@ Appendix C Additional configurations that satisfy the integrated KdV equation (12)
The precise QA configuration with larger negative shear in Fig. 21 was obtained by a series of optimizations starting from the precise QA.Landreman and Paul (2022) The mean shear $\bar{\hat{s}}$ was extracted from the precise QA configuration, and a new least-squares term $(\bar{\hat{s}}-\bar{\hat{s}}_{*})^{2}$ was added to the objective function, where the target value for mean shear $\bar{\hat{s}}_{*}$ is taken to be slightly different from the extracted value. This new objective is then optimized locally, using the precise QA as the initial state. This ensures that the additional least-squares term is almost zero, so the precise QA is approximately an optimum to the modified objective function. Hence, the new local optimum can be expected to be close to the initial configuration. This process is then repeated starting from the slightly modified optimum, each step generating a similar equilibrium with a slightly different mean shear.
Mean shear is here defined as $b/\bar{\iota}$, where $b$ is the result of fitting a first-order polynomial $a+bs$ to the iota profile, where $s$ is the normalized toroidal flux; $\bar{\iota}$ is the rotational transform averaged over all radii. The mean shear in the original precise QA is about 0.018-0.018, and we rounded this to 0.02-0.02 incremented the target shear of each optimization in steps of $0.005$ to generate configurations with a range of shear. Fig. 21 and Fig. 22 show example with negative and positive shear, respectively.
@@
We show equilibria in Figures 23 and 24 that were not derived from the precise QA/QH. Instead, they were only optimized for quasisymmetry on the outermost flux surface, with the quasisymmetry degrading as the axis is approached. For both cases, the ratio of symmetry-breaking modes to quasisymmetric modes near the axis is $\sim 10^{-3}$. While this may seem a small value, the resulting ripple creates enough noise in the $(dB/d\ell)^{2}$ vs. $B$ plot to render it meaningless, and the error for all fits is $\sim 1$ in that region. The symmetry-breaking modes were filtered out when calculating the roots shown in panels (c). Note that the equilibrium shown in Fig. 23 is the same one used in Fig. 2 in the main text.